To save for their retirement, a couple deposits $4000 in an account that pays 5.9% annual interest compounded quarterly. What will be the value of their investment after 20 years?
Answer by nerdybill(7384)

To save for their retirement, a couple deposits $4000 in an account that pays 5.9% annual interest compounded quarterly. What will be the value of their investment after 20 years?
.
A = P(1 + r/n)^(nt)
where
A is amount after time t
P is the initial amount
r is the interest rate
n is the number of time it is compounded
.
Plugging in what is given:
A = 4000(1 + .059/4)^(4*20)
A = 4000(1 + .059/4)^80
A = 4000(1 + 0.01475)^80
A = 4000(1.01475)^80
A = 4000(3.226)
A = $12905.80